# Electronic correlations arising from anti-Stoner spin excitations: an ab initio study of itinerant ferro- and antiferromagnet

### v1

@@ Abstract
The anti-Stoner excitations are a spin-flips in which, effectively, an electron is promoted from a minority to a majority spin state, i.e., complementary to Stoner excitations and spin-waves. Since their spectral power is negligible in strong itinerant ferromagnets and they are identically absent in the ferromagnetic Heisenberg model, their properties and role in correlating electrons were hardly investigated so far. On the other hand, they are present in weak ferromagnets, fcc Ni being a prominent example, and both types of spin-flips (down-to-up and up-to-down) must be treated on the equal footing in systems with the degenerate spin up and down bands, in particular antiferromagnets in which case we choose CrSb as a model system. For these two materials we evaluate the strength of the effective interaction between the quasiparticles and the gas of virtual spin-flip excitations. To this end, we compute the corresponding self-energy taking advantage of our novel efficient ab initio numerical scheme [1]. We find that in Ni the band-structure renormalization due to the anti-Stoner processes is weaker than the one due to Stoner-type magnons in the majority spin channel but the two become comparable in the minority one. The effect can be traced back primarily to the spectral strength of the respective spin excitations and the densities of the final available quasiparticle states in the scattering process. For the antiferromagnet, the situation is more complex and we observe that the electron-magnon interaction is sensitive not only to these densities of states but critically to the spatial shapes of the coupling magnonic modes as well.
@@ I Introduction
Antiferromagnets (AFMs) were predicted long ago as magnetic materials containing two magnetic sublattices with antiparallel orientation of their magnetic moments and zero net magnetization [2]; now this term is frequently applied to any magnetically ordered state with zero net magnetization. Actually, there are much more AFMs than ferromagnets in nature, and the study of AFMs is a very important part of the modern physics of magnetism [3, 4, 5, 6]. The absence of stray fields and relatively weak sensitivity to external magnetic fields in comparison with ferro- and ferrimagnets can be a serious advantage for the emerging fields of spintronics and magnonics [7, 8]. Furthermore, their ultrafast spin dynamics grants spintronic device operation in the terahertz frequency range [9]. Among the recent experimentally realized milestones in spintronic and magnonic using AFMs are the detection of ultrafast generation of spin currents [10], the coherent spin wave transport including optical excitation and detection [11] and long distance magnon transport in layered van der Waals AFMs [12]. Furthermore, several unexpected phenomena like anomalous Hall effects, unconventional transport and the emergence of skyrmions which are connected to the topological behavior of the material were recently observed [13, 14, 15].
@@
One of the main theoretcal tools utilized to study correlated electrons is the dynamic mean field theory (DMFT) often used in combination with DFT [28, 29, 30]. Several studies of AFMs using DMFT can be found in the literature [31, 32, 33]. The main drawback of the DMFT lies in the local approximation which causes a wave-vector independent self energy. One can go beyond DMFT using e.g. dual fermion method considering DMFT solution as zero-order approximation [34, 35, 18]. Alternative approaches using many body perturbation theory (MBPT), often in the formulation of Hedin [36], are able to account for non-locality, but are often only applied to very simple materials due to their numerical complexity [37, 38]. However, we recently proposed a method based on MBPT which is able to account for the non-locality of the self energy while still being numerically feasible for complex materials [1]. This was achieved by approximating complex quantities of MBPT with their counterparts in the theory of time dependent DFT [39]. While the formalism presented in [1] is only applicable to strong ferromagnets, it can easily be expanded to being able to account for all thinkable collinear magnetic structures as will be shown in this work. In principle, this generalization also allows to consider paramagnets, but this aspect goes beyond the scope of the current report.
@@ II.2 Linear response time-dependent density functional theory
Linear response TDDFT allows to compute the retarded counterpart of the reducible density-density response function which can be used to extract the spectral functions. It is achieved in the two step procedure. First, the formally non-interacting Kohn-Sham response function is computed. It is given by
where the sum includes Kohn-Sham states with energy and state vector while represents the Fermi-Dirac statistic and is the th Pauli matrix. Greek indices represent spin directions. For the spin-flip channels addressed this work, the transverse response is obtained by setting .
@@ III.1 Anti-Stoner processes in fcc nickel
The plots show that the anti-Stoner excitations are of primarily single-particle character and the action of the in the susceptibility Dyson equation does not lead to the formation of collective modes or sharp features in the channel. This is in striking contrast to the formation of spin-waves from the interacting Stoner excitations in its counterpart. Furthermore, the spectral power contained in the , as evident from Fig. 2, is clearly smaller than in the Stoner case. However, as will be shown below, the impact of the quasiparticle scattering of these excitations leads to a sizable renormalization of the electronic band structure. Thus, the contribution to the self-energy must be systematically incorporated in the description of itinerant ferromagnets.
We proceed now with the evaluation of the self-energy without the = 0 approximation. due to the lack of well-defined collective modes in the -channel, no “magnon-pole”-like approximation can be constructed here, as it is done for the spectral function [1]. Instead, is taken directly from the linear response TDDFT. In order to account for the double counting [45], the Kohn-Sham part of the correlation function must be subtracted from its enhanced counterpart, yielding
@@ III.2 Antiferromagnet CrSb
We consider the layered AFM CrSb in the hexagonal NiAs structure. Due to the high Néel temperature of this material has been identified as particularly promising in the field of AFM-spintronics. Furthermore, recent studies have shown that thin films of CrSb are ferrimagnetic or can even feature the altermagnetic behavior [49, 50]. Here, we investigate the impact of electron-magnon coupling on the electronic properties of its bulk phase.
@@
The chirality of magnon excitations in AFM has been already carefully discussed before [42]. The magnons break the symmetry between up and down sublattices. The up-to-down spin-flip is associated with the up sublattice precessing stronger (with larger amplitude) than the down one. Conversely, the down-to-up spin-flip causes the down sublattice to precess stronger than the up one. In both cases the magnon causes the system to feature a net magnetic moment which is absent in the ground state. These intricate symmetries are reflected in the properties of the self-energy discussed below.
@@
Here, the subscript labels the spin of the quasi-particle while the superscripts represent the site-resolved blocks of the self-energy, i.e., Cr site with upwards or downwards magnetization (see also equation 4). These symmetries are caused by the fact that the electronic states for both spin directions are degenerate and the magnon spectrum obeys the symmetries discussed above. Consequently, in antiferromagnets the electronic spectrum of both spin channels are renormalized in the same manner upon the action of the magnon-electron interaction. In the specific case of CrSb, the energy of electronic bands changes weakly, cf. lower panel in Fig. 5. However, they experience rather strong damping caused by the scattering with the virtual magnons. Note that the damping of the electronic states is strongly dependent on the momentum of the state.
@@
First of all, contrary to the case of strong FM systems, both up and down electrons and holes can involve in a sizable interaction with magnons. A further difference is manifested by the fact that in the FM systems the strength of the scattering is primarily determined by the availability of the final quasiparticle states while in for the antiferromagnets the spatial shape of the magnons modes facilitating the decay is of paramount importance as well.
@@
On the other hand, the up holes on the Cr⇑ site can emit magnons and decay efficiently, is large. Their decay process requires a Stoner-like excitations of the Fermi liquid which feature large amplitudes on the Cr⇑ sites. While the density of final down holes is limited on the site, it is still sizable. The intensity of decay process of the up electrons on the Cr⇓ sites can be qualitatively explained using this same argument. Of course, the carriers can also decay into states on other Cr sites, the processes being described by the off-diagonal parts of the self-energy and . These dominating quasiparticle-magnon scattering processes are depicted schematically in Fig. 7.
With the self-energy at hand, the renormalized Fermi level and the magnetic moments can be computed. Due to the high numerical costs, this is done in a one-shot procedure, i.e., not self-consistently. It turns out that the Fermi energy decreases by approximately 0.2 eV. The magnetic moment of the Cr atoms is reduced from 2.84 (LSDA approximation) to 2.67 (LSDA+ approximation including the magnon-electron scattering) which agrees with available experimental results [55]. The renormalized density of states is shown in the upper panel in Fig. 6.
We conclude by discussing briefly the role of the non-magnetic antimony atoms in the magnon-electron scattering. When resorting to the magnon-pole approximation, their contribution vanishes identically, as they carry no magnetic moment in the ground state. Furthermore, it means that the DOS of the Sb sites is unchanged in this approximation. However, in general the non-magnetic moments could become magnetically polarized when neighboring magnetic moments cease to be collinear during a magnon excitation and the approximation might break. Nevertheless, the magnetic polarizability of Sb in the CrSb system is minuscule which we confirmed by computing the full magnetic susceptibility from linear response TDDFT for selected wave-vectors. On the other hand, all atoms, including the Sb, are fully included in the description of the electronic structure given by the KKR Green’s function.
@@ IV Summary
In this report, we completed the ab initio description of magnon-electron scattering processes in collinear magnets. On the example of the fcc Ni, we investigated carefully the correlations emerging from the interaction of quasiparticles and the gas of virtual anti-Stoner excitations. While its influence is weaker than the one of Stoner-type magnons, our results show that the strong ferromagnet approximation does not hold universally for magnon-induced electron damping in all itinerant ferromagnets. In the particular case of fcc Ni, the impact of Stoner- and anti-Stoner-like exciations on the quasiparticle spectra becomes of comaprable strength in the minority spin channel while in the majority one the Stoner-type excitations dominate.
In the antiferromagnetic systems, as shown on the example of CrSb, the electron-magnon interaction is sensitive not only to the density of the available final particle states in the scattering processes but to the spatial shape of coupling magnonic modes as well. This result is in striking contrast to the ferromagnetic materials like nickel for which the density of final states is the main factor impacting the strength of the quasiparticle renormalization. Furthermore, the antiferromagnetic symmetries of the electronic structure and chirality of the magnon excitations lead to the quasiparticle damping which acts selectively based on the atomic site and its magnetic moment. We believe it to offer a new approach to carrier life-time engineering in spintronic devices.

### v2

@@ Abstract
The anti-Stoner excitations are a spin-flips in which, effectively, an electron is promoted from a minority to a majority spin state, i.e., complementary to Stoner excitations and spin-waves. Since their spectral power is negligible in strong itinerant ferromagnets and they are identically absent in the ferromagnetic Heisenberg model, their properties and role in correlating electrons were hardly investigated so far. On the other hand, they are present in weak ferromagnets, fcc Ni being a prominent example, and both types of spin-flips (down-to-up and up-to-down) must be treated on equal footing in systems with degenerate spin up and down bands, in particular antiferromagnets in which case we choose CrSb as a model system. For these two materials we evaluate the strength of the effective interaction between the quasiparticles and the gas of virtual spin-flip excitations. To this end, we compute the corresponding self-energy taking advantage of our novel efficient ab initio numerical scheme [1]. We find that in Ni the band-structure renormalization due to the anti-Stoner processes is weaker than the one due to Stoner-type magnons in the majority spin channel but the two become comparable in the minority one. The effect can be traced back primarily to the spectral strength of the respective spin excitations and the densities of the final available quasiparticle states in the scattering process. For the antiferromagnet, the situation is more complex and we observe that the electron-magnon interaction is sensitive not only to these densities of states but critically to the spatial shapes of the coupling magnonic modes as well.
@@ I Introduction
Antiferromagnets (AFMs) were predicted long ago as magnetic materials containing two magnetic sublattices with antiparallel orientation of their magnetic moments and zero net magnetization [2]; now this term is frequently applied to any magnetically ordered state with zero net magnetization. Actually, there are much more AFMs than ferromagnets in nature, and the study of AFMs is a very important part of the modern physics of magnetism [3, 4, 5, 6]. The absence of stray fields and relatively weak sensitivity to external magnetic fields in comparison with ferro- and ferrimagnets can be a serious advantage for the emerging fields of spintronics and magnonics [7, 8]. Furthermore, their ultrafast spin dynamics grants spintronic device operation in the terahertz frequency range [9]. Among the recent experimentally realized milestones in spintronic and magnonic using AFMs are the detection of ultrafast generation of spin currents [10], the coherent spin wave transport including optical excitation and detection [11] and long distance magnon transport in layered van der Waals AFMs [12]. Furthermore, several unexpected phenomena like anomalous Hall effects, unconventional transport and the emergence of skyrmions were recently observed [13, 14, 15].
@@
One of the main theoretical tools utilized to study correlated electrons is the dynamic mean field theory (DMFT) often used in combination with DFT [28, 29, 30]. Several studies of AFMs using DMFT can be found in the literature [31, 32, 33]. The main drawback of the DMFT lies in the local approximation which causes a wave-vector independent self energy. One can go beyond DMFT using e.g. the dual fermion method considering the DMFT solution as zero-order approximation [34, 35, 18]. Alternative approaches using many body perturbation theory (MBPT), often in the formulation of Hedin [36], are able to account for non-locality, but are often only applied to very simple materials due to their numerical complexity [37, 38]. However, we recently proposed a method based on MBPT which is able to account for the non-locality of the self energy while still being numerically feasible for complex materials [1]. This was achieved by approximating complex quantities of MBPT with their counterparts in the theory of time dependent DFT [39]. While the formalism presented in [1] is only applicable to strong ferromagnets, it can easily be expanded to being able to account for all thinkable collinear magnetic structures as will be shown in this work. In principle, this generalization also allows to consider paramagnets, but this aspect goes beyond the scope of the current report.
@@ II.2 Linear response time-dependent density functional theory
Linear response TDDFT allows to compute the retarded counterpart of the reducible density-density response function which can be used to extract the spectral functions. First, the formally non-interacting Kohn-Sham response function is computed. It is given by
where the sum includes Kohn-Sham states with energy and state vector while represents the Fermi-Dirac statistic and is the th Pauli matrix. Greek indices represent spin directions. For the spin-flip channels addressed in this work, the transverse response is obtained by setting .
@@ III.1 Anti-Stoner processes in fcc nickel
The plots show that the anti-Stoner excitations () are of primarily single-particle character in both Kohn-Sham and interacting system, i.e. no collective modes or sharp features are present. This is in striking contrast to the formation of spin-waves from the interacting Stoner excitations in its counterpart. Furthermore, the spectral power contained in , as evident from Fig. 2, is clearly smaller than in the Stoner case. However, as will be shown below, the impact of the quasiparticle scattering of these excitations leads to a sizable renormalization of the electronic band structure. Thus, the contribution to the self-energy must be systematically incorporated in the description of itinerant ferromagnets.
We proceed now with the evaluation of the self-energy without the = 0 approximation. due to the lack of well-defined collective modes in the -channel, no “magnon-pole”-like approximation can be constructed here, as it is done for the spectral function [1]. Instead, is taken directly from the linear response TDDFT. In order to account for the double counting issue [45], the Kohn-Sham part of the correlation function must be subtracted from its enhanced counterpart, yielding
@@ III.2 Antiferromagnet CrSb
We consider the layered AFM CrSb in the hexagonal NiAs structure. Due to its high Néel temperature this material has been identified as particularly promising in the field of AFM-spintronics. Furthermore, recent studies have shown that thin films of CrSb are ferrimagnetic or can even feature altermagnetic behavior [49, 50]. Here, we investigate the impact of electron-magnon coupling on the electronic properties of its bulk phase.
@@
The chirality of magnon excitations in AFM has already been carefully discussed before [42]. The magnons break the symmetry between the sublattices with upwards () and downwards () magnetization. The up-to-down spin-flip is associated with the up sublattice precessing stronger (with larger amplitude) than the down one. Conversely, the down-to-up spin-flip causes the down sublattice to precess stronger than the up one. In both cases the magnon causes the system to feature a net magnetic moment which is absent in the ground state. These intricate symmetries are reflected in the properties of the self-energy discussed below.
@@
Here, the subscript labels the spin of the quasi-particle while the superscripts represent the site-resolved blocks of the self-energy, i.e., Cr site with upwards or downwards magnetization (see also equation 4). These symmetries are caused by the fact that the electronic states for both spin directions are degenerate and the magnon spectrum obeys the symmetries discussed above. Consequently, in antiferromagnets the electronic spectra of both spin channels are renormalized in the same manner upon the action of the magnon-electron interaction. In the specific case of CrSb, the energy of electronic bands changes weakly, cf. lower panel in Fig. 5. However, they experience rather strong damping caused by the scattering with the virtual magnons. Note that the damping of the electronic states is strongly dependent on the momentum of the state.
@@
First of all, contrary to the case of strong FM systems, both up and down electrons and holes can involve in a sizable interaction with magnons. A further difference is manifested by the fact that in the FM systems the strength of the scattering is primarily determined by the availability of the final quasiparticle states while for antiferromagnets the spatial shape of the magnon modes facilitating the decay is of paramount importance as well.
@@
On the other hand, the up holes on the Cr⇑ site can emit magnons and decay efficiently, is large. Their decay process requires Stoner-like excitations of the Fermi liquid which feature large amplitudes on the Cr⇑ sites. While the density of final down holes is limited on the site, it is still sizable. The intensity of the decay processes of the up electrons on the Cr⇓ sites can be qualitatively explained using this same argument. Of course, the carriers can also decay into states on other Cr sites, the processes being described by the off-diagonal parts of the self-energy and . These dominating quasiparticle-magnon scattering processes are depicted schematically in Fig. 7.
With the self-energy at hand, the renormalized Fermi level and the magnetic moments can be computed. Due to the high numerical cost, this is done in a one-shot procedure, i.e., not self-consistently. It turns out that the Fermi energy decreases by approximately 0.2 eV. The magnetic moment of the Cr atoms is reduced from 2.84 (LSDA approximation) to 2.67 (LSDA+ approximation including the magnon-electron scattering) which agrees with available experimental results [55]. The renormalized density of states is shown in the upper panel in Fig. 6.
We conclude by briefly discussing the role of the non-magnetic antimony atoms in the magnon-electron scattering. When resorting to the magnon-pole approximation, their contribution vanishes identically, as they carry no magnetic moment in the ground state. Furthermore, it means that the DOS of the Sb sites is unchanged in this approximation. However, in general the non-magnetic atoms could become magnetically polarized when neighboring magnetic moments cease to be collinear during a magnon excitation and the approximation might break. Nevertheless, the magnetic polarizability of Sb in the CrSb system is minuscule which we confirmed by computing the full magnetic susceptibility from linear response TDDFT for selected wave-vectors. On the other hand, all atoms, including the Sb, are fully included in the description of the electronic structure given by the KKR Green’s function.
@@ IV Summary
In this report, we completed the ab initio description of magnon-electron scattering processes in collinear magnets. On the example of fcc Ni, we investigated carefully the correlations emerging from the interaction of quasiparticles and the gas of virtual anti-Stoner excitations. While its influence is weaker than the one of Stoner-type magnons, our results show that the strong ferromagnet approximation does not hold universally for magnon-induced electron damping in all itinerant ferromagnets. In the particular case of fcc Ni, the impact of Stoner- and anti-Stoner-like exciations on the quasiparticle spectra becomes of comaprable strength in the minority spin channel while in the majority one the Stoner-type excitations dominate.
In antiferromagnetic systems, as shown on the example of CrSb, the electron-magnon interaction is sensitive not only to the density of available final particle states in the scattering processes, but to the spatial shape of coupling magnonic modes as well. This result is in striking contrast to ferromagnetic materials like nickel for which the density of final states is the main factor impacting the strength of the quasiparticle renormalization. Furthermore, the antiferromagnetic symmetries of the electronic structure and chirality of the magnon excitations lead to the quasiparticle damping which acts selectively based on the atomic site and its magnetic moment. We believe it to offer a new approach to carrier life-time engineering in spintronic devices.